calculate the radius of gyration of a hoop of radius 5 centimetre rotating about an Axis passing through its centre and normal to its plane

Grade:11

2 Answers

Arun
4 years ago
The radius of gyration of a hoop of  radius 5.00 cm rotating about an axis passing through its centre and normal to its plane.
I(tangent) = 2/5 Ma² + MR²
∴  I(tangent) = 7/5 MR²
∴ Mk² = 7/5 MR²
∴ k = √(7/5 x 5² = √35 cm
 
Vikas TU
4 years ago
Dear student 
The radius of gyration of hoop of Radius 5.00cm rotating about an axis passing through its centre and normal to its plane 
I(tangent) = 2/5 Ma^2 + MR^2 
MK^2 = 7/5 MR^2 
K^2 = 7/5 R^2 
k = sqrt(35) 
Hope this information will clear your doubts about the topic .
If you have any more doubts just ask here on the forum and our experts will try to help you out as soon as possible